# Data-driven Software-based Power Estimation for Embedded Devices

### v1

@@ III-C2 Decision Tree Model
The following two models use the decision tree model. The accuracy of 2 decision tree models is shown in Table II.
TABLE II: The accuracy of decision tree models
@@ IV-A5 Portability
This module offers a robust solution for monitoring CPU performance, enhancing the understanding of power consumption and enabling optimized resource management across diverse computing environments. The Data Tracker module’s design and implementation reflect a significant contribution to performance analysis tools, providing a reliable and versatile mechanism for real-time CPU monitoring.
@@ IV-B Example: Power Prediction for programs with Multiple Frequencies
Figure 7 shows a portion of the real trace derived from the Data Tracker module. The trace includes the utilization and frequency of a workload (top) and the predicted power values (bottom) for each frequency segment. The trace is the experimental result for the benchmark $basicmath$ running with variable frequencies under the $ondemand$ governor. The x-axis represents time, while the y-axis conveys information about each frequency segment. In the upper part of Figure 7, the height of the orange bar indicates the frequency level, and the height of the blue bar denotes the utilization at that frequency. The lower part of Figure 7 shows the predicted power values (depicted by the red dashed line) for each frequency segment, calculated using our constructed power model.
Figure 7: A real trace of utilization and frequency and it’s prediction
@@ V Evaluation of Power Prediction and Discussion
In this section, we rigorously evaluate our models by comparing the predicted results with empirical data for programs executed under variable frequencies and discuss the implications of our findings. The evaluation includes testing programs that runs under the Linux $ondemand$ governor, which dynamically adjusts frequencies based on system utilization. We also deploy our approach on a different embedded platform to validate its reproducibility.
@@ V-A Evaluation Under $ondemand$ Governor
Figure 8 shows the comparison between the measured data and the predicted data for 11 benchmarks. The red bars represent the measured data, while the blue bars indicate the data predicted by our model. As observed, the predicted and actual values are closely aligned, signifying the high prediction accuracy of our model.
Figure 8: Benchmark predictions vs real values on JTN
Table III presents the absolute error for each benchmark’s prediction. The mean absolute error across all benchmarks is 0.1118, and the mean squared error is 0.0164. These metrics confirm that the prediction error is minimal, thereby validating the high prediction accuracy of our model.
TABLE III: Absolute error
@@ V-B Reproducibility on Different Platforms
Figure 9 presents the benchmark predictions versus real values on the Raspberry Pi. The results indicate that the average mean error is approximately 4%, with the best test achieving an average error of 2%. These results demonstrate that the methodology is applicable to different embedded platforms for achieving accurate energy consumption predictions.
Figure 9: Benchmark predictions vs real values on RBP
However, certain discrepancies were observed. For instance, the $fileio$ benchmark behaves differently on the Raspberry Pi. This is due to the specific behavior of file I/O operations on this platform.
@@
Most small inexpensive embedded devices in MEC systems do not have software-based power measurement support. This paper proposes a smart data-driven method using consumer-grade meters to construct software-level power estimation models for such embedded devices. We first describe the two drawbacks associated with the use of a power meter to measure energy consumption, which are the inability to measure instantaneous power consumption and the inability to act on energy feedback-based systems. Then, we describe our approach to constructing the power model by systematically generating workloads and collecting training data. To use our approach, one can follow the steps described in this paper to build the model for their own devices. The steps include building the workload, collecting training data, constructing the model and using the model. During the experiments, we explored several types of models, and in the end, the model that worked best on the device that we tested was a per-frequency polynomial regression model on utilization. The model accuracy is 92% compared to the long-duration measurement. The approach is simple and systematic, capable of achieving good results in predicting energy consumption, thus allowing the researchers/application developers to avoid using a power meter for power management. We also provide a platform-independent kernel module that can be used to collect running traces of utilization and frequencies needed for power prediction for programs running in Linux environment.

### v2

@@ III-C2 Decision Tree Model
The following two models use the decision tree model. The accuracy of 2 decision tree models is shown in Table V.
TABLE V: The accuracy of decision tree models
@@ IV-A5 Portability
The Data Tracker module’s design and implementation contribute to performance analysis, provide a reliable and versatile mechanism for real-time CPU monitoring, enhance the understanding of power consumption, and enable optimized resource management across diverse computing environments.
@@ IV-B Example: Power Prediction for programs with Multiple Frequencies
Fig. 7 shows a portion of the real trace derived from the Data Tracker module. The trace includes the utilization and frequency of a workload (top-left) and the predicted power values and Energy Consumption (bottom-left) for each frequency segment. The trace is the experimental result for the benchmark $basicmath$ running with variable frequencies under the $ondemand$ governor. The x-axis represents time, while the y-axis conveys information about each frequency segment. In the upper part of Fig. 7, the height of the striped blue bar indicates the frequency level, and the height of the red square markers and lines denotes the utilization at that frequency. The lower part of Fig. 7 shows the predicted power values (depicted by the red circular markers and lines) for each frequency segment, calculated using our constructed power model, and the energy consumption calculated by duration times power (depicted by the height of the striped blue bar). The right part of Fig. 7 depicts how the power of each segment is calculated (top to bottom) and the calculation across time (left to right).
Figure 7: Actual trace of utilization and frequency with predicted power consumption
@@ V Evaluation of Power Prediction and Discussion
In this section, we evaluate our models by comparing the predicted results with empirical data for programs executed under variable frequencies and discuss the implications of our findings. The evaluation includes testing programs that runs under the Linux $ondemand$ governor, which dynamically adjusts frequencies based on system utilization. We also deploy our approach on a different embedded platform to validate its reproducibility.
@@ V-A Evaluation Under $ondemand$ Governor
Fig. 8 shows the comparison between the measured data and the predicted data for 11 benchmarks. The dark blue bars represent the measured values, while the striped light blue bars indicate the values predicted by our model. As observed, the predicted and actual values are closely aligned, signifying the high prediction accuracy of our model.
Figure 8: Benchmark predictions vs. real values on JTN
Table VI presents the absolute error for each benchmark’s prediction. The mean absolute error across all benchmarks is 0.1118, and the mean squared error is 0.0164. These metrics confirm that the prediction error is minimal, thereby validating the high prediction accuracy of our model.
TABLE VI: Absolute error
@@ V-B Reproducibility on Different Platforms
Fig. 9 presents the benchmark predictions versus real values on the Raspberry Pi. The results indicate that the average mean error is approximately 4%, with the best test achieving an average error of 2%. These results demonstrate that the methodology is applicable to different embedded platforms for achieving accurate energy consumption predictions.
Figure 9: Benchmark predictions vs. real values on RBP
However, certain discrepancies were observed. For instance, the $fileio$ benchmark behaves differently on the Raspberry Pi. This could be due to the specific behavior of file I/O operations on this platform.
V-C Limitations
While our proposed method demonstrates high accuracy in power estimation for devices like the Jetson Nano and Raspberry Pi, it has certain limitations.
Our power estimation model focuses on applications with primarily CPU-Centric power consumption. That is, the CPU is the primary energy consumer in embedded devices. For applications running on devices where other components such as GPUs, high-speed memory (e.g., LPDDR4 RAM), or solid-state drives (SSDs) have significant power consumption, our model may not accurately capture the total energy usage. For instance, devices performing intensive graphics processing with GPUs, conducting high-throughput memory operations, or engaging in frequent disk I/O with SSDs may exhibit power profiles our CPU-focused model does not account for, leading to reduced overall accuracy in power estimation for such scenarios.
Our method has been tested on devices with homogeneous multi-core CPUs, such as the Jetson Nano and Raspberry Pi, which consist of identical cores. The applicability of our method to heterogeneous systems with different types of cores, such as big.LITTLE architectures commonly found in modern mobile processors, has not been evaluated. In these systems, cores may have different performance and power characteristics, making power estimation more complex.
@@ V-C Limitations
Most small inexpensive embedded devices in MEC systems do not have software-based power measurement support. This paper proposes a smart data-driven method using consumer-grade meters to construct software-level power estimation models for such embedded devices. We first describe the drawbacks associated with the use of a power meter to measure energy consumption, which are the inability to measure instantaneous power consumption and the inability to act on energy feedback-based systems. Then, we describe our approach to constructing the power model by systematically generating workloads and collecting training data. To use our approach, one can follow the steps described in this paper to build the model for their own devices. The steps include building the workload, collecting training data, constructing the model and using the model. During the experiments, we explored several types of models, and in the end, the model that worked best on the device that we tested was a per-frequency polynomial regression model on utilization. The model accuracy is 92% compared to the long-duration measurement. The approach is straightforward and systematic, capable of achieving good results in predicting energy consumption. This allows the researchers/application developers to eliminate the need of using a power meter for power measurement once the model is established. We also provide a platform-independent kernel module that can be used to collect running traces of utilization and frequencies needed for power prediction for programs running in Linux environment.